# Generative Flows on Synthetic Pathway for Drug Design

### v1

@@ 1 Introduction
While generative models have shown success in molecular discovery with desirable biological properties, most overlook synthesizability which is a crucial factor for wet-lab validation (Gao & Coley, 2020). One line to improve synthesizability is multi-objective optimization using cheap functions to estimate the synthesizability (Ertl & Schuffenhauer, 2009), but this is too simplified to reflect complex synthetic principles (Cretu et al., 2024). Other efforts aim to project molecules from generative models into a synthesizable space (Gao et al., 2022b; Luo et al., 2024), but chemical modifications in this process can degrade the optimized properties.
@@ 2 Related Works
Structure-based drug discovery. The first type of SBDD involves pocket-specific optimization methods to enhance docking scores against a single pocket, including evolutionary algorithms (Reidenbach, 2024), reinforcement learning (RL) (Zhavoronkov et al., 2019), and GFlowNets (Bengio et al., 2021). However, these require individual optimizations for each pocket, limiting scalability. The second type is based on pocket-conditional generation, which generates molecules against arbitrary given pockets without additional training. This can be achieved by distribution-based generative models (Ragoza et al., 2022; Peng et al., 2022; Guan et al., 2023b; Schneuing et al., 2023) trained on protein-ligand complex datasets to model ligand distributions for given pockets. On the other hand, Shen et al. (2023) formulated a pocket-conditioned policy for GFlowNets to generate samples from reward-biased distributions in a zero-shot manner.
Syntheis-oriented generative models. To ensure the synthesizability of generated molecules, synthesis-oriented de novo design approaches incorporate combinatorial chemistry principles into generative models. Bradshaw et al. (2019) represented the synthetic pathways as directed acyclic graphs (DAG) for generative modeling. Horwood & Noutahi (2020) formulated the synthetic pathway generation as an MDP and optimize the molecules with RL. Similarly, Gao et al. (2022b) employed a genetic algorithm to optimize synthesis trees to generate molecules with the desired properties. Seo et al. (2023) proposed a conditional generative model to directly sample molecules with desired properties without optimization. Recently, Cretu et al. (2024); Koziarski et al. (2024) trained the models with the GFlowNets objective to generate diverse and potent molecules.
@@ 3.3 Flow Network on Action Space Subsampling
However, large action spaces require numerous parameters which increase model complexity. To address this, we use an additional network $\phi^{\text{block}}_{\theta}$ for AddFirstReactant and ReactBi, which embeds the building block $b$ into a continuous action space with its structural information:
@@
GFlowNet training. In this work, we use the trajectory balance (TB; Malkin et al., 2022) as the training objective of GFlowNets from Eq. 9 and train models following Algorithm 1. The action space subsampling is performed for each transition $s_{t}\rightarrow s_{t+1}$: $\mathcal{A}^{\ast}_{t}\sim\mathcal{P}(\mathcal{A})$.
@@
Algorithm 1 Training GFlowNets with action space subsampling
1:input Entire action space $\mathcal{A}$, Maximum trajectory length $N$
2:repeat
6:until model converges
@@ 4 Experiments
Overview. We validate the effectiveness of RxnFlow in two common SBDD tasks: pocket-specific optimization (Sec. 4.1) and pocket-conditional generation (Sec. 4.2). To the best of our knowledge, this is the first synthesis-oriented approach for pocket-conditional generation. We also investigate the applicability of RxnFlow in real-world drug discovery pipelines where new further objectives may be introduced (Sec. 4.3) and the building block libraries are constantly expanded (Sec. 4.4). Lastly, we conduct an ablation study in Sec. 4.5 and a theoretical analysis in Sec. D.3.
@@ 4.1 Pocket-Specific Optimization with GPU-accelerated Docking
Percentage of Synthesizable Molecules (%, $\uparrow$)
@@
Setup. Since GFlowNets sample a large number of molecules for online training, we employed a GPU-accelerated UniDock (Yu et al., 2023) with Vina scoring (Trott & Olson, 2010). It is well known that docking can be hacked by increasing molecule size (Pan et al., 2003), so the appropriate constraints are required. We select QED (Bickerton et al., 2012) as a comprehensive molecular property constraint, QED$>$0.5, and set the reward function as $R(x)=w_{1}\text{QED}(x)+w_{2}\widehat{\text{Vina}}(x)$ where $w_{1},w_{2}$ are used as the input of multi-objective GFlowNets (Jain et al., 2023) and are set to 0.5 for non-GFlowNet baselines. $\widehat{\text{Vina}}$ is a normalized docking score (Eq. 30).
@@
Each method generates up to 64,000 molecules for each of the 15 proteins in the LIT-PCBA dataset (Tran-Nguyen et al., 2020). We then filter the molecules with the property constraint and select the top 100 diverse candidates based on the docking score, using a Tanimoto distance threshold of 0.5 to ensure structural diversity. The selected molecules are evaluated with the following metrics: Hit ratio (%) measures the fraction of hits, defined as the molecules that are identified as synthesizable by AiZynthFinder and having better docking scores than known active ligands (Lee et al., 2023). Vina (kcal/mol) measures the average docking score. Synthesizability (%) is the fraction of synthesizable molecules. Synthetic complexity, which is highly correlated to yield and cost, is evaluated as the average number of synthesis steps (Coley et al., 2018).
@@
Baselines. We perform comparisons to various synthetic-oriented approaches: genetic algorithm (SynNet) (Gao et al., 2022b), conditional generative model (BBAR [^1] Since BBAR requires labeled training data with QED and docking score, we perform docking with random 62,720 ZINC molecules for training and evaluate 1,280 samples according to the reported splitting ratio.) (Seo et al., 2023), and GFlowNets (SynFlowNet, RGFN) (Cretu et al., 2024; Koziarski et al., 2024). For SynFlowNet and RGFN, we used 6,000 and 350 blocks, respectively, and set the maximum reaction step to 4 following the original papers. Moreover, we consider fragment-based GFlowNets (FragGFN) to analyze the effects of synthetic constraints on the performance. For FragGFN, we also consider additional synthesizability objectives with commonly-used synthetic accessibility score (SA; Ertl & Schuffenhauer, 2009) (FragGFN+SA).
@@
Results. The results for the first five targets are shown in Tables 4 and 4, and additional results for the 10 remaining targets are reported in Sec. D.1. RxnFlow outperforms the baselines across all test proteins, demonstrating that the expanded sample space with the large action space enabled the model to generate more potent and diverse molecules. Additionally, as shown in Tables 4 and 4, RxnFlow ensures the synthesizability of the generated molecules more effectively than the other synthesis-oriented methods and GFlowNets which employ the same reactions as ours. These results support our primary assertion that existing synthesis-oriented approaches using more synthetic steps with smaller building block libraries can increase overall synthesis complexity and reduce synthesizability. Furthermore, FragGFN+SA does not show meaningful improvement in synthesizability, implying that optimization of a cheap synthesizability estimation function is suboptimal.
@@ 4.2 Zero-Shot Sampling via Pocket Conditioning
Table 5: CrossDocked2020 benchmark. We report the average and median values over the average properties for each test pocket. The best results are in bold and the second ones are in underlined. We denote the Validity as Valid, Synthesizability as Synthesiz, and Diversity as Div. Reference means the known binding active molecules in the test set.
Valid ($\uparrow$) Vina ($\downarrow$) QED ($\uparrow$) Synthesiz ($\uparrow$) Div ($\uparrow$) Time ($\downarrow$)
@@
Atom Pocket2Mol 098.3% -7.60 -7.16 0.57 0.58 29.1% 22.0% 0.83 2504a
TargetDiff 091.5% -7.37 -7.56 0.49 0.49 09.9% 03.2% 0.87 3428a
DecompDiff 066.0% -8.35 -8.25 0.37 0.35 00.9% 00.0% 0.84 6189a
Fragment TacoGFN 100.0% -8.24 -8.44 0.67 0.67 01.3% 01.0% 0.67 0004b
Reaction RxnFlow 100.0% -8.85 -9.03 0.67 0.67 34.8% 34.5% 0.81 0004b
@@
Setup. We extend our works to a pocket-conditional generation problem to design binders for arbitrary pockets without additional training oracles (Ragoza et al., 2022; Liu et al., 2022; Peng et al., 2022; Guan et al., 2023a; b; Schneuing et al., 2023). To address this challenge, we follow TacoGFN (Shen et al., 2023), which is a fragment-based GFlowNet for pocket-conditional generation. Since it requires more training oracles to learn pocket-conditional policies than target-specific generation, TacoGFN used pre-trained proxies that predict docking scores against arbitrary pockets using pharmacophore representation (Seo & Kim, 2023). Since RxnFlow explicitly considers synthesizability, we exclude the SA score from the TacoGFN’s reward function as described in Sec. C.2.
We generate 100 molecules for each of the 100 test pockets in the CrossDocked2020 benchmark (Francoeur et al., 2020) and evaluate them with the following metrics: Vina (kcal/mol) measures the average docking score from QuickVina (Alhossary et al., 2015). QED measures the average drug-likeness of molecules. Synthesizability (%) is the fraction of synthesizable molecules. Diversity measures an average pairwise Tanimoto distance of ECFP4 fingerprints (Morgan, 1965). Moreover, we report the Validity (%) which is the percentage of unique RDKit-readable molecules, and Time (sec.) which is the average sampling time to generate 100 molecules.
Baselines. We compare RxnFlow with state-of-the-art distribution learning-based models trained on a synthesizable drug set: an autoregressive model (Pocket2Mol (Peng et al., 2022)) and diffusion models (TargetDiff (Guan et al., 2023a), DecompDiff (Guan et al., 2023b)). We also perform a comparison with an optimization-based TacoGFN (Shen et al., 2023). For a fair comparison with the distribution learning-based approaches, we used the docking proxy trained on CrossDocked2020.
Result. As shown in Table 5, RxnFlow achieves significant improvements in drug-related properties. In particular, RxnFlow attains the best average docking score of -8.85 kcal/mol, significantly outperforming TacoGFN at -8.24 kcal/mol and the state-of-the-art diffusion model (DecompDiff) at -8.35 kcal/mol while achieving high drug-like properties. Moreover, RxnFlow ensures the synthesizability comparable to known active ligands, outperforming the fragment-based TacoGFN trained on the SA score objective and the distributional learning-based models trained on synthesizable drug molecules. Figure 3 illustrates generated molecules against TANK-binding kinase 1 (TBK1) which is not included in the training set.
@@ 4.3 Introducing Additional Objective without Retraining
In drug discovery, new objectives often arise during the research process, such as enhancing solubility, reducing toxicity, or improving selectivity (Fink et al., 2022; Joshi et al., 2021). These additional objectives typically not only require retraining models but also increase the optimization complexity. In this context, RxnFlow can achieve some additional objectives by simply introducing constraints to MDP without retraining thanks to the non-hierarchical action space (Sec. 3.4).
As shown in Figure 5, we explore the scenario of adding a solubility objective to a pre-trained GFlowNet in Sec. 4.2. Specifically, we target the generation of hydrophobic molecules by restricting the building blocks with topological polar surface area (TPSA) in the bottom 15% (“low”) and sampled 500 molecules for the KRAS-G12C mutant (PDB Id: 6oim). While there are slight differences in the QED distributions due to the correlation between TPSA and QED, the generated molecules are more hydrophobic and retain similar overall reward distributions.
@@ 4.5 Ablation Study: the effect of building block library size
Expanding the size of building block libraries provides an opportunity to discover more diverse and potent drug candidates (Grygorenko et al., 2020). In Sec. 4.1, RxnFlow outperforms SynFlowNet and RGFN which use smaller block libraries, but differences in model architectures may have contributed to these results. To isolate the effect of the building block library size, we conduct an ablation study using partial libraries with a pocket-specific optimization task against the kappa-opioid receptor (PDB Id: 6b73), as illustrated in Figure 6(a-c). The results indicate that increasing the library size enhances both optimization power, in terms of docking scores, and diversity, in terms of a higher number of unique Bemis-Murcko scaffolds (Bemis & Murcko, 1996) and an increased Tanimoto diversity of the generated molecules. Additionally, as shown in Figure 6(d), the generation time only doubles on the 100-fold larger action space, highlighting the efficiency of RxnFlow. These results demonstrate the forte of RxnFlow in navigating a broader chemical space to discover novel drug candidates by overcoming the computational limitations for expanding the action space.

### v2

@@ 1 Introduction
While generative models have shown success in molecular discovery with desirable biological properties, most overlook synthesizability which is a crucial factor for wet-lab validation (Gao & Coley, 2020). One line to improve synthesizability is multi-objective optimization using cheap functions to estimate the synthesizability (Ertl & Schuffenhauer, 2009), but this is too simplified to reflect complex synthetic principles (Cretu et al., 2024). Other efforts aim to project molecules from generative models into a synthesizable space (Gao et al., 2022b; Luo et al., 2024; Gao et al., 2024b), but chemical modifications in this process can degrade the optimized properties.
@@ 2 Related Works
Structure-based drug discovery. The first type of SBDD involves pocket-specific optimization methods to enhance docking scores against a single pocket, including evolutionary algorithms (Reidenbach, 2024), reinforcement learning (RL) (Zhavoronkov et al., 2019), and GFlowNets (Bengio et al., 2021). However, these require individual optimizations for each pocket, limiting scalability. The second type is based on pocket-conditional generation, which generates molecules against arbitrary given pockets without additional training. This can be achieved by distribution-based generative models (Ragoza et al., 2022; Peng et al., 2022; Guan et al., 2023b; Schneuing et al., 2023; Qu et al., 2024) trained on protein-ligand complex datasets to model ligand distributions for given pockets. On the other hand, Shen et al. (2023) formulated a pocket-conditioned policy for GFlowNets to generate samples from reward-biased distributions in a zero-shot manner.
Syntheis-oriented generative models. To ensure the synthesizability of generated molecules, synthesis-oriented de novo design approaches incorporate combinatorial chemistry principles into generative models. Bradshaw et al. (2019) represented the synthetic pathways as directed acyclic graphs (DAG) for generative modeling. Horwood & Noutahi (2020) formulated the synthetic pathway generation as an MDP and optimized the molecules with RL. Similarly, Gao et al. (2022b) employed a genetic algorithm to optimize synthesis trees to generate molecules with the desired properties. Seo et al. (2023) proposed a conditional generative model to directly sample molecules with desired properties without optimization. Recently, Cretu et al. (2024); Koziarski et al. (2024) proposed the reaction-based GFlowNet to generate diverse and potent molecules.
@@ 3.3 Flow Network on Action Space Subsampling
However, large action spaces require numerous parameters which increase model complexity. To address this, we use an additional network $\phi^{\text{block}}_{\theta}$ for AddFirstReactant and ReactBi, which embeds the building block $b$ into a continuous action space with its structural information, molecular fingerprints (see Sec. B.3):
@@
GFlowNet training. In this work, we use the trajectory balance (TB; Malkin et al., 2022) as the training objective of GFlowNets from Eq. 9 and train models following Sec. B.4. The action space subsampling is performed for each transition $s_{t}\rightarrow s_{t+1}$: $\mathcal{A}^{\ast}_{t}\sim\mathcal{P}(\mathcal{A})$.
@@ 4 Experiments
Overview. We validate the effectiveness of RxnFlow in two common SBDD tasks: pocket-specific optimization (Sec. 4.1) and pocket-conditional generation (Sec. 4.2). To the best of our knowledge, this is the first synthesis-oriented approach for pocket-conditional generation. We also investigate the applicability of RxnFlow in real-world drug discovery pipelines where new further objectives may be introduced (Sec. 4.3) and the building block libraries are constantly expanded (Sec. 4.4). Lastly, we conduct an ablation study in Sec. 4.5 and a theoretical analysis in Sec. D.8. Source code available at https://github.com/SeonghwanSeo/RxnFlow.
@@ 4.1 Pocket-Specific Optimization with GPU-accelerated Docking
AiZynthFinder Success Rate (%, $\uparrow$)
@@
Setup. Since GFlowNets sample a large number of molecules for online training, we employed a GPU-accelerated UniDock (Yu et al., 2023) with Vina scoring (Trott & Olson, 2010). It is well known that docking can be hacked by increasing molecule size (Pan et al., 2003), so the appropriate constraints are required. We select QED (Bickerton et al., 2012) as a comprehensive molecular property constraint, QED$>$0.5, and set the reward function as $R(x)=w_{1}\text{QED}(x)+w_{2}\widehat{\text{Vina}}(x)$ where $w_{1},w_{2}$ are used as the input of multi-objective GFlowNets (Jain et al., 2023) for all GFlowNets and are set to 0.5 for non-GFlowNet baselines. $\widehat{\text{Vina}}$ is a normalized docking score (Eq. 32).
@@
Each method generates up to 64,000 molecules for each of the 15 proteins in the LIT-PCBA dataset (Tran-Nguyen et al., 2020). We then filter the molecules with the property constraint and select the top 100 diverse candidates based on the docking score, using a Tanimoto distance threshold of 0.5 to ensure structural diversity. The selected molecules are evaluated with the following metrics: Hit ratio (%) measures the fraction of hits, defined as the molecules that are identified as synthesizable by AiZynthFinder and having better docking scores than known reference ligands (Lee et al., 2023). Vina (kcal/mol) measures the average docking score. Synthesizability (%) is the fraction of synthesizable molecules identified by AiZynthFinder. Synthetic complexity, which is highly correlated to yield and cost, is evaluated as the average number of synthesis steps (Coley et al., 2018).
@@
Baselines. We perform comparisons to various synthetic-oriented approaches: genetic algorithm (SynNet) (Gao et al., 2022b), conditional generative model (BBAR) (Seo et al., 2023), and GFlowNets (SynFlowNet, RGFN) (Cretu et al., 2024; Koziarski et al., 2024). For SynFlowNet and RGFN, we used 6,000 and 350 blocks, respectively, and set the maximum reaction step to 4 following the original papers. Moreover, we consider fragment-based GFlowNets (FragGFN) to analyze the effects of synthetic constraints on the performance. For FragGFN, we also consider additional synthesizability objectives with commonly-used synthetic accessibility score (SA; Ertl & Schuffenhauer, 2009) (FragGFN+SA).
@@
Results. The results for the first five targets are shown in Tables 4 and 4, and additional results for the 10 remaining targets are reported in Sec. D.1. The property distribution for each target are reported in Sec. D.2. RxnFlow outperforms the baselines across all test proteins, demonstrating that the expanded sample space with the large action space enabled the model to generate more potent and diverse molecules. Additionally, as shown in Tables 4 and 4, RxnFlow ensures the synthesizability of the generated molecules more effectively than the other synthesis-oriented methods and GFlowNets which employ the same reactions as ours. These results support our primary assertion that existing synthesis-oriented approaches using more synthetic steps with smaller building block libraries can increase overall synthesis complexity and reduce synthesizability. Furthermore, FragGFN+SA does not show meaningful improvement in synthesizability, implying that optimization of a cheap synthesizability estimation function is suboptimal.
@@ 4.2 Zero-Shot Sampling via Pocket Conditioning
Table 5: CrossDocked2020 benchmark. We report the average and median values over the average properties for each test pocket. The best results are in bold and the second ones are in underlined. We denote the Generation Success as Succ., Synthesizability as Synthesiz., and Diversity as Div. Reference means the known binding active molecules in the test set. MolCRAFT-large (Qu et al., 2024) is the result when generating with more atoms than the reference ligands.
Succ. ($\uparrow$) Vina ($\downarrow$) QED ($\uparrow$) Synthesiz. ($\uparrow$) Div. ($\uparrow$) Time ($\downarrow$)
@@
Atom Pocket2Mol 098.3% -7.60 -7.16 0.57 0.58 29.1% 22.0% 0.83 02504
TargetDiff 091.5% -7.37 -7.56 0.49 0.49 09.9% 03.2% 0.87 03428
DecompDiff 066.0% -8.35 -8.25 0.37 0.35 00.9% 00.0% 0.84 06189
DiffSBDD 076.0% -6.95 -7.10 0.47 0.48 02.9% 02.0% 0.88 00135
MolCRAFT 096.7% -8.05 -8.05 0.50 0.50 16.5% 09.1% 0.84 00141
MolCRAFT-large 070.8% -9.25 -9.24 0.45 0.44 03.9% 00.0% 0.82 $>$141
Fragment TacoGFN 100.0% -8.24 -8.44 0.67 0.67 01.3% 01.0% 0.67 00004
Reaction RxnFlow 100.0% -8.85 -9.03 0.67 0.67 34.8% 34.5% 0.81 00004
@@
Setup. We extend our works to a pocket-conditional generation problem to design binders for arbitrary pockets without additional training oracles (Ragoza et al., 2022; Liu et al., 2022; Peng et al., 2022; Guan et al., 2023a; b; Schneuing et al., 2023). To address this challenge, we follow TacoGFN (Shen et al., 2023), which is a fragment-based GFlowNet for pocket-conditional generation. Since it requires more training oracles to learn pocket-conditional policies than target-specific generation, TacoGFN used pre-trained proxies that predict docking scores against arbitrary pockets using pharmacophore representation (Seo & Kim, 2024). Since RxnFlow explicitly considers synthesizability, we exclude the SA score from the TacoGFN’s reward function as described in Sec. C.2.
We generate 100 molecules for each of the 100 test pockets in the CrossDocked2020 benchmark (Francoeur et al., 2020) and evaluate them with the following metrics: Vina (kcal/mol) measures the average docking score from QuickVina (Alhossary et al., 2015). QED measures the average drug-likeness of molecules. Synthesizability (%) is the fraction of synthesizable molecules. Diversity measures an average pairwise Tanimoto distance of ECFP4 fingerprints. Moreover, we report the Generation Success (%) which is the percentage of unique RDKit-readable molecules without disconnected parts, and Time (sec.) which is the average sampling time to generate 100 molecules.
Baselines. We compare RxnFlow with state-of-the-art distribution learning-based models trained on a synthesizable drug set: an autoregressive model (Pocket2Mol (Peng et al., 2022)), diffusion models (TargetDiff (Guan et al., 2023a), DiffSBDD (Schneuing et al., 2023), DecompDiff (Guan et al., 2023b)), and bayesian flow network (MolCRAFT (Qu et al., 2024)). We also perform a comparison with an optimization-based TacoGFN (Shen et al., 2023). For a fair comparison with the distribution learning-based approaches, we used the docking proxy trained on CrossDocked2020.
Result. As shown in Table 5, RxnFlow achieves significant improvements in drug-related properties. In particular, RxnFlow outperforms the docking score for TacoGFN and attains high drug-likeness while showing a competitive docking score with the state-of-the-art model, MolCRAFT-large. Moreover, RxnFlow ensures the synthesizability comparable to known active ligands, outperforming the fragment-based TacoGFN trained on the SA score objective and the distributional learning-based models trained on synthesizable drug molecules. Figure 3 illustrates generated molecules against TANK-binding kinase 1 (TBK1) which is not included in the training set.
@@ 4.3 Introducing Additional Objective without Retraining
In drug discovery, new objectives often arise during the research process, such as enhancing solubility, reducing toxicity, or improving selectivity (Fink et al., 2022; Joshi et al., 2021). These additional objectives typically not only require retraining models but also increase the optimization complexity. In this context, RxnFlow can achieve some additional objectives by simply introducing constraints to MDP without retraining thanks to the non-hierarchical action space (Sec. 3.4). As shown in Figure 5, we explore the scenario of adding a solubility objective to a pre-trained GFlowNet in Sec. 4.2. Specifically, we target the generation of hydrophobic molecules by restricting the building blocks with topological polar surface area (TPSA) in the bottom 15% (“low”) and sampled 500 molecules for the KRAS-G12C mutant (PDB Id: 6oim). While there are slight differences in the QED distributions due to the correlation between TPSA and QED, the generated molecules are more hydrophobic and retain similar overall reward distributions. We also performed the ablation study of non-hierarchical MDP in Sec. D.6.
@@ 4.5 Ablation Study: the effect of building block library size
Expanding the size of building block libraries provides an opportunity to discover more diverse and potent drug candidates (Grygorenko et al., 2020). In Sec. 4.1, RxnFlow outperforms SynFlowNet and RGFN which use smaller block libraries, but differences in model architectures may have contributed to these results. To isolate the effect of the building block library size, we conduct an ablation study using partial libraries with a pocket-specific optimization task against the kappa-opioid receptor (PDB Id: 6b73), as illustrated in Figure 6(a-c). The results indicate that increasing the library size enhances both optimization power, in terms of docking scores, and diversity, in terms of a higher number of unique Bemis-Murcko scaffolds (Bemis & Murcko, 1996) and an increased Tanimoto diversity of the generated molecules. Additionally, as shown in Figure 6(d), the generation time only doubles on the 100-fold larger action space, highlighting the efficiency of RxnFlow. These results demonstrate the forte of RxnFlow in navigating a broader chemical space to discover novel drug candidates by overcoming the computational limitations for expanding the action space. We also investigated the scaling laws with other reaction-based GFlowNets in Sec. D.3.
Figure 7: Affect of 3D interaction modeling. Mean and standard error over 3 seeds
4.6 Towards Further Development
Our framework has room for improvement regarding more efficient learning and sampling. First, the 3D interaction modeling can enhance the generative performance. Given the high correlation between binding affinity and conformation, such spatial relationships could provide meaningful information for our generative model to make a reasonable decision. In Figure 7, we observed that the 3D interaction modeling using docking conformations boosts the discovery of candidate molecules against the beta-2-adrenergic receptor. Second, the current action space subsampling method forces exploration due to uniform sampling to minimize bias. We can enhance the exploitation by prioritizing the building blocks of action space subsampling instead of uniform subsampling.